OpenStudy (anonymous):

What is the equation of the line that passes through the point (1, 6) and has a slope of 3?

OpenStudy (anonymous):

Start with the point slope formula: point (h,k), slope m \[y-k = m(x-h)\]

OpenStudy (anonymous):

in your case the point (h,k) = (1,6)

OpenStudy (anonymous):

and the slope m = 3

OpenStudy (anonymous):

y-6=3(1-k)

OpenStudy (anonymous):

Close! but you mixed up the m(x-h) part.

OpenStudy (anonymous):

you got the m plugged in right, but not the h, and you lost your x

OpenStudy (anonymous):

The way i was taught was y- y1=m(x-x1

OpenStudy (anonymous):

y-6=3(x-1)

OpenStudy (anonymous):

Yes, that's right. I just call y1 k, and x1 h cause some people find that less confusing.

OpenStudy (anonymous):

y-6=3x-3 Distributed

OpenStudy (anonymous):

Yep. Then add 6 to both sides.

OpenStudy (anonymous):

y-6=3x-3 +6 +6 y=3x-3

OpenStudy (anonymous):

y= 3x-3

OpenStudy (anonymous):

mixed up the sign. it should be y = 3x + 3

OpenStudy (anonymous):

I have another question im having trouble with

OpenStudy (anonymous):

x - 3 > -9 and 3x 15

OpenStudy (anonymous):

3x <= 15

OpenStudy (anonymous):

Ok first one first.

OpenStudy (anonymous):

add 3 to both sides of the inequality.

OpenStudy (anonymous):

x >-6

OpenStudy (anonymous):

3x <= 15 --- --- 3 3 x <= 5

OpenStudy (anonymous):

x <= 5

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

On a number line which way would they travel and open circle or close circles?